A cylinder has a radius of 2.1 cm and a length of 8.8 cm. Total charge 6.1*10-7 C is distributed uniformly throughout. The volume charge density is:
1
Expert's answer
2020-07-30T11:00:53-0400

p = Charge Volume Density

"p= \frac{q}{v}"

"q = 6.1 \times10^{-7} C"

"v = \pi r^{2} h"

"v = \pi\times 2.1\times 2.1\times8.8\times 10^{-2}"

"v = 122.23\times10^{-6} m^{3}"

"p = \frac{6.1\times 10^{-7}}{122.23\times 10^{-6} }"

"p = 0.04\times10^{-6} C/m^{3}"